A man spent one fifth of his money at the butchers and one third at the market. He had £5.60 left. How much did he start with


1
Expert's answer
2021-05-21T07:22:47-0400

let the amount be x

"butchers=1\/5x"

"market=1\/3x"

total spent"=\frac{8x}{15}"

"remainder=7x\/15=5.6"

"x=\frac{15\times 5.6}{7}=" £ "12"